# Adversarial Robust Decision Transformer: Enhancing Robustness of RvS via Minimax Returns-to-go

### v1

@@ Abstract
Decision Transformer (DT), as one of the representative Reinforcement Learning via Supervised Learning (RvS) methods, has achieved strong performance in offline learning tasks by leveraging the powerful Transformer architecture for sequential decision-making. However, in adversarial environments, these methods can be non-robust, since the return is dependent on the strategies of both the decision-maker and adversary. Training a probabilistic model conditioned on observed return to predict action can fail to generalize, as the trajectories that achieve a return in the dataset might have done so due to a weak and suboptimal behavior adversary. To address this, we propose a worst-case-aware RvS algorithm, the Adversarial Robust Decision Transformer (ARDT), which learns and conditions the policy on in-sample minimax returns-to-go. ARDT aligns the target return with the worst-case return learned through minimax expectile regression, thereby enhancing robustness against powerful test-time adversaries. In experiments conducted on sequential games with full data coverage, ARDT can generate a maximin (Nash Equilibrium) strategy, the solution with the largest adversarial robustness. In large-scale sequential games and continuous adversarial RL environments with partial data coverage, ARDT demonstrates significantly superior robustness to powerful test-time adversaries and attains higher worst-case returns compared to contemporary DT methods.
@@ 1 Introduction
Reinforcement Learning via Supervised Learning (RvS) (Emmons et al., 2021), has garnered attention within the domain of offline Reinforcement Learning (RL). Framing RL as a problem of outcome-conditioned sequence modeling, these methodologies have showcased strong performance in offline benchmark tasks (Wu et al., 2024; Yamagata et al., 2023; Janner et al., 2021; Schmidhuber, 2019; Kumar et al., 2019). One of the representative RvS methods is Decision Transformer (DT)  (Chen et al., 2021), which simply trains a policy conditioned on target return via behavior cloning loss. Given the apparent efficacy and simplicity of RvS, our attention is drawn to exploring its performance in adversarial settings. In this paper, we aim to use RvS to achieve adversarial robustness, a critical capability for RL agents to manage environmental variations and adversarial disturbances (Pinto et al., 2017; Tessler et al., 2019; Kamalaruban et al., 2020; Vinitsky et al., 2020; Curi et al., 2021; Rigter et al., 2022). Such challenges are prevalent in real-world scenarios, e.g., changes in road conditions and decision-making among multiple autonomous vehicles.
In offline RL with adversarial actions, the naive training of action-prediction models conditioned on the history and observed returns such as DT, or expected returns such as ESPER and DoC (Paster et al., 2022; Yang et al., 2022; Yamagata et al., 2023), may result in a non-robust policy. This issue arises from the potential distributional shifts in the policy of the adversary during offline learning. Specifically, if the behavior policies used for data collection are suboptimal, a well-trained conditional policy overfitted to the training dataset may fail to achieve the desired target return during testing, if the adversary’s policy at test time has changed or become more effective.
@@
Main Contributions. Consequently, there is a need for data curation to indicate the potential worst-case returns associated with each action in the Decision Transformer (DT), improving the robustness against adversarial actions. Building on this concept, we introduce the first robust training algorithm for Decision Transformer, the Adversarial Robust Decision Transformer (ARDT). To the best of our knowledge, this paper represents the first exploration of the robustness of RvS methods from an adversarial perspective. In ARDT, we leverage Expectile Regression  (Newey and Powell, 1987; Aigner et al., 1976) to transform the original returns-to-go to in-sample minimax returns-to-go, and train DT with the relabeled returns. In the experiments, we illustrate the robustness of ARDT in three settings (i) Games with full data coverage (ii) A long-horizon discrete game with partial data coverage (iii) Realistic continuous adversarial RL problems. We provide evidence showcasing the robustness of ARDT to more powerful adversary compared to the behavior policy and the distributional shifts of the adversary’s policy in the test time. Furthermore, in the adversarial MuJoCo tasks, namely Noisy Action Robust MDP  (Tessler et al., 2019), ARDT exhibits better robustness to a population of adversarial perturbations compared to existing DT methods.
Algorithms Policy ($\hat{R}=6$): (${P}(a_{0})$, ${P}(a_{1})$)
DT (0.5400, 0.4600)
ERC-DT (0.9998, 0.0002)
ARDT (ours) (0.0006, 0.9994)
@@
Related work. Reinforcement Learning via Supervised Learning is a promising direction for solving general offline RL problems by converting the RL problem into the outcome-conditioned prediction problem (Kumar et al., 2019; Schmidhuber, 2019; Ghosh et al., 2019; Paster et al., 2020; Emmons et al., 2021). Decision Transformer (DT) is one of the RvS instance (Chen et al., 2021; Kim et al., 2023), leveraging the strength of the sequence model Transformer (Vaswani et al., 2017). The variants of DT have been extensively investigated to address the trajectory stitching (Wu et al., 2024; Yamagata et al., 2023), stochasticity (Paster et al., 2020; Ortega et al., 2021; Paster et al., 2022; Yang et al., 2022), goal-conditioned problems (Furuta et al., 2021), and generalization in different environments (Meng et al., 2021; Xu et al., 2022; Lee et al., 2022; Xu et al., 2023). For different problems, it is critical to select an appropriate target as the condition of the policy. $Q$-Learning DT (Yamagata et al., 2023) relabels the trajectories with the estimated $Q$-values learned to achieve trajectory stitching (Kumar et al., 2020). ESPER (Paster et al., 2022), aims at solving the environmental stochasticity via relabeling the trajectories with expected returns-to-go. Although our method also transform the returns-to-go to estimated values, we instead associate worst-case returns-to-go with trajectories, aiming at improving its robustness against adversarial perturbations.
Adversarial RL can be framed as a sequential game; with online game-solving explored extensively in literature (Lanctot et al., 2009; Heinrich et al., 2015; Sunehag et al., 2017; Lowe et al., 2017; Lanctot et al., 2017; Srinivasan et al., 2018; Schrittwieser et al., 2020; Tang et al., 2023). In the offline setting, where there is no interaction with the environment, solving games is challenging due to distributional shifts in the adversary’s policy, necessitating adequate data coverage. Tabular methods, such as Nash $Q$-Learning—a decades-old approach for decision-making in competitive games (Hu and Wellman, 2003)—and other similar methods (Cui and Du, 2022; Zhong et al., 2022) address these challenges through pessimistic out-of-sample estimation. While value-based methods are viable for addressing adversarial problems, our focus is on enhancing the robustness of RvS methods in these settings. Note that while some offline multi-agent RL methods, including DT-based approaches, are designed for cooperative games (Meng et al., 2021; Wen et al., 2022; Tseng et al., 2022; Wang et al., 2024), our research is oriented towards competitive games.
@@ 2 Problem Setup
We consider the adversarial reinforcement learning problem involving a protagonist and an adversary, as described by Pinto et al. (2017). This setting is rooted in the two-player Zero-Sum Markov Game framework Littman (1994); Perolat et al. (2015). This game is formally defined by the tuple $(\mathcal{S},\mathcal{A},\bar{\mathcal{A}},T,R,p_{0})$. Here, $\mathcal{S}$ represents the state space, where $\mathcal{A}$ and $\bar{\mathcal{A}}$ denote the action spaces for the protagonist and adversary, respectively. The reward function $R:\mathcal{S}\times\mathcal{A}\times\bar{\mathcal{A}}\to\mathbb{R}$ and the transition kernel $T:\mathcal{S}\times\mathcal{A}\times\bar{\mathcal{A}}\to\Delta(\mathcal{S})$ depend on the state and the joint actions of both players. The initial state distribution is denoted by $p_{0}$. At each time-step $t\leq H$, both players observe the state $s_{t}\in\mathcal{S}$ and take actions $a_{t}\in\mathcal{A}$ and $\bar{a}_{t}\in\bar{\mathcal{A}}$. The adversary is considered adaptive if its action $\bar{a}_{t}$ depends on $s_{t}$ and $a_{t}$. The protagonist receives a reward $r_{t}=R(s_{t},a_{t},\bar{a}_{t})$, while the adversary receives a negative reward $-r_{t}$.
@@ 2.1 RvS in Adversarial Environments
We extend the RvS framework by incorporating adversarial settings within the Hindsight Information Matching (HIM) paradigm (Furuta et al., 2021). In adversarial HIM, the protagonist receives a goal (denoted by $z$) and acts within an adversarial environment. The protagonist’s performance is evaluated based on how well it achieves the presented goal. Mathematically, the protagonist aims to minimize a distance metric $D(\cdot,\cdot)$ between the target goal $z$ and the information function $I(\cdot)$ applied to a trajectory $\tau$. This sub-trajectory follows a distribution $\rho$ dependent on both the protagonist’s policy $\pi_{z}=\pi(a_{t}\mid\tau_{0:t-1},s_{t},z)$ and the test-time adversarial policy $\bar{\pi}_{\textnormal{test}}$. Formally, the information distance is minimized as follows:
@@
where $\rho^{\pi_{z},\bar{\pi}_{\textnormal{test}}}$ represents the trajectory distribution obtained by executing rollouts in a Markov Game with the protagonist’s policy $\pi_{z}$ and the test-time adversarial policy $\bar{\pi}_{\textnormal{test}}$. Under HIM framework, Decision Transformer Chen et al. (2021) uses adopt return-to-go value as the information function, while ESPER Paster et al. (2022) employs the expected return-to-go. The well-trained RvS policy, including DT and ESPER, is derived from behavior trajectories filtered based on the condition $I(\tau)=z$:
@@
Let $\pi_{\mathcal{D}}$ and $\bar{\pi}_{\mathcal{D}}$ be the data collecting policies used to gather data for training an RvS protagonist policy $\pi$. Assume $T(s_{t+1}\mid s_{t},a_{t},{\bar{a}_{t}})=\rho^{\pi_{\mathcal{D}},\bar{\pi}_{\mathcal{D}}}(s_{t+1}\mid\tau_{0:t-1},s_{t},a_{t},{\bar{a}_{t}},I(\tau)=z)$  [^1] This condition can be readily met in environments formulated as a game with deterministic transitions, which are prevalent in many sequential games (e.g., Go and Chess (Schrittwieser et al., 2020)), as well as in frequently considered MuJoCo environments involving adversarial action noise (Pinto et al., 2017; Cheng et al., 2022). , for any goal $z$ such that $\rho^{\pi_{\mathcal{D}},\bar{\pi}_{\mathcal{D}}}(I(\tau)=z\mid s_{0})>0$, the information distance can be minimized: $\mathbb{E}_{\tau\sim\rho^{\pi_{z},\bar{\pi}_{\textnormal{test}}}}\!\left[{D(I(\tau_{t}),z)}\right]=0$ if $\bar{\pi}_{\textnormal{test}}(\bar{a}_{t}\mid\tau_{0:t-1},s_{t},a_{t})=\rho^{\pi_{\mathcal{D}},\bar{\pi}_{\mathcal{D}}}(\bar{a}_{t}\mid\tau_{0:t-1},s_{t},a_{t},I(\tau)=z)$.
@@
Figure 2: Training of Adversarial Robust Decision Transformer. We adopt Expectile Regression for estimator $\widetilde{Q}$ to approximate the in-sample minimax. In the subsequent protagonist DT training, we replace the original returns-to-go with the learned values $\widetilde{Q}^{*}$ to train policy.
3 Adversarial Robust Decision Transformer
To tackle the problem of suboptimal adversarial behavior in the dataset, we propose Adversarial Robust Decision Transformer (ARDT), a worst-case-aware RvS algorithm. In ARDT, the information function is defined as the expected minimax return-to-go:
@@ 3 Adversarial Robust Decision Transformer
Intuitively, this information function represents the expected return when an adversary aims to minimize the value, while the protagonist subsequently seeks to maximize it. To train a protagonist policy as described in Eq. (3), ARDT first relabels the trajectory with the minimax returns-to-go in Eq. (4). Secondly, the protagonist policy is then trained similarly to a standard Decision Transformer (DT). This two-step overview of ARDT is provided in Figure 2, where the left block includes the minimax return estiation with expectile regression for trajectory relabeling, and the right block represents the DT training conditioned on the relabeled returns-to-go.
@@
We adopt Expectile Regression (ER) (Newey and Powell, 1987; Aigner et al., 1976) to approximate $\widetilde{Q}$. This choice is particularly advantageous for RL applications because it helps avoid out-of-sample estimation or the need to query actions to approximate the minimum or maximum return (Kostrikov et al., 2021). Specifically, the loss function for ER is a weighted mean squared error (MSE):
@@
Algorithm 1 Adversarial Robust Decision Transformer (ARDT)
@@ 4 Experiments
In this section, we conduct experiments to examine the robustness of our algorithm, Adversarial Robust Decision Transformer (ARDT), in three settings: (i) Short-horizon sequential games, where the offline dataset has full coverage and the test-time adversary is optimal (Section 4.1), (ii) A long-horizon sequential game, Connect Four, where the offline dataset has only partial coverage and the distributional-shifted test-time adversary (Section 4.2), and (iii) The standard continuous Mujoco tasks in the adversarial setting and a population of test-time adversaries (Section 4.3). We compare our algorithm with baselines including Decision Transformer and ESPER. The implementation deatils are in Appendix C. Notably, since the rewards and states tokens encompass sufficient information about the adversary’s actions, all DT models are implemented not conditioned on the past adversarial tokens to reduce the computational cost.

### v2

# Adversarially Robust Decision Transformer

@@ Abstract
Decision Transformer (DT), as one of the representative Reinforcement Learning via Supervised Learning (RvS) methods, has achieved strong performance in offline learning tasks by leveraging the powerful Transformer architecture for sequential decision-making. However, in adversarial environments, these methods can be non-robust, since the return is dependent on the strategies of both the decision-maker and adversary. Training a probabilistic model conditioned on observed return to predict action can fail to generalize, as the trajectories that achieve a return in the dataset might have done so due to a suboptimal behavior adversary. To address this, we propose a worst-case-aware RvS algorithm, the Adversarially Robust Decision Transformer (ARDT), which learns and conditions the policy on in-sample minimax returns-to-go. ARDT aligns the target return with the worst-case return learned through minimax expectile regression, thereby enhancing robustness against powerful test-time adversaries. In experiments conducted on sequential games with full data coverage, ARDT can generate a maximin (Nash Equilibrium) strategy, the solution with the largest adversarial robustness. In large-scale sequential games and continuous adversarial RL environments with partial data coverage, ARDT demonstrates significantly superior robustness to powerful test-time adversaries and attains higher worst-case returns compared to contemporary DT methods.
@@ 1 Introduction
Reinforcement Learning via Supervised Learning (RvS) [8], has garnered attention within the domain of offline Reinforcement Learning (RL). Framing RL as a problem of outcome-conditioned sequence modeling, these methodologies have showcased strong performance in offline benchmark tasks [50, 53, 13, 37, 17]. One of the representative RvS methods is Decision Transformer (DT)  [2], which simply trains a policy conditioned on target return via behavior cloning loss. Given the apparent efficacy and simplicity of RvS, our attention is drawn to exploring its performance in adversarial settings. In this paper, we aim to use RvS to achieve adversarial robustness, a critical capability for RL agents to manage environmental variations and adversarial disturbances [34, 42, 14, 46, 5, 36]. Such challenges are prevalent in real-world scenarios, e.g., changes in road conditions and decision-making among multiple autonomous vehicles.
In offline RL with adversarial actions, the naive training of action-prediction models conditioned on the history and observed returns such as DT, or expected returns such as ESPER and DoC [32, 54, 53], which we refer to as Expected Return-Conditioned DT (ERC-DT), may result in a non-robust policy. This issue arises from the potential distributional shifts in the policy of the adversary during offline learning. Specifically, if the behavior policies used for data collection are suboptimal, a well-trained conditional policy overfitted to the training dataset may fail to achieve the desired target return during testing, if the adversary’s policy at test time has changed or become more effective.
@@
Main Contributions. Consequently, there is a need for data curation to indicate the potential worst-case returns associated with each action in the Decision Transformer (DT), improving the robustness against adversarial actions. Building on this concept, we introduce the first robust training algorithm for Decision Transformer, the Adversarially Robust Decision Transformer (ARDT). To the best of our knowledge, this paper represents the first exploration of the robustness of RvS methods from an adversarial perspective. In ARDT, we leverage Expectile Regression  [29, 1] to transform the original returns-to-go to in-sample minimax returns-to-go, and train DT with the relabeled returns. In the experiments, we illustrate the robustness of ARDT in three settings (i) Games with full data coverage (ii) A long-horizon discrete game with partial data coverage (iii) Realistic continuous adversarial RL problems. We provide evidence showcasing the robustness of ARDT to more powerful adversary compared to the behavior policy and the distributional shifts of the adversary’s policy in the test time. Furthermore, in the adversarial MuJoCo tasks, namely Noisy Action Robust MDP  [42], ARDT exhibits better robustness to a population of adversarial perturbations compared to existing DT methods.
Algorithms Policy ($\hat{R}=6$): ${P}(a_{0})$, ${P}(a_{1})$
DT 0.54, 0.46
ERC-DT 0.9998, 0.0002
ARDT (ours) 0.0006, 0.9994
@@
Related work. Reinforcement Learning via Supervised Learning is a promising direction for solving general offline RL problems by converting the RL problem into the outcome-conditioned prediction problem [17, 37, 10, 31, 8]. Decision Transformer (DT) is one of the RvS instance [2, 15], leveraging the strength of the sequence model Transformer [45]. The variants of DT have been extensively investigated to address the trajectory stitching [50, 53], stochasticity [31, 30, 32, 54], goal-conditioned problems [9], and generalization in different environments [26, 51, 21, 52]. For different problems, it is critical to select an appropriate target as the condition of the policy. $Q$-Learning DT [53] relabels the trajectories with the estimated $Q$-values learned to achieve trajectory stitching [18]. ESPER [32], aims at solving the environmental stochasticity via relabeling the trajectories with expected returns-to-go. Although our method also transform the returns-to-go to estimated values, we instead associate worst-case returns-to-go with trajectories, aiming at improving its robustness against adversarial perturbations.
Adversarial RL can be framed as a sequential game; with online game-solving explored extensively in literature [19, 11, 40, 24, 20, 39, 38, 41]. In the offline setting, where there is no interaction with the environment, solving games is challenging due to distributional shifts in the adversary’s policy, necessitating adequate data coverage. Tabular methods, such as Nash $Q$-Learning—a decades-old approach for decision-making in competitive games [12]—and other similar methods [4, 56] address these challenges through pessimistic out-of-sample estimation. While value-based methods are viable for addressing adversarial problems, our focus is on enhancing the robustness of RvS methods in these settings. Note that while some offline multi-agent RL methods, including DT-based approaches, are designed for cooperative games [26, 48, 44, 47], our research is oriented towards competitive games.
@@ 2 Problem Setup
We consider the adversarial reinforcement learning problem involving a protagonist and an adversary, as described by Pinto et al. [34]. This setting is rooted in the two-player Zero-Sum Markov Game framework [23, 33]. This game is formally defined by the tuple $(\mathcal{S},\mathcal{A},\bar{\mathcal{A}},T,R,p_{0})$. Here, $\mathcal{S}$ represents the state space, where $\mathcal{A}$ and $\bar{\mathcal{A}}$ denote the action spaces for the protagonist and adversary, respectively. The reward function $R:\mathcal{S}\times\mathcal{A}\times\bar{\mathcal{A}}\to\mathbb{R}$ and the transition kernel $T:\mathcal{S}\times\mathcal{A}\times\bar{\mathcal{A}}\to\Delta(\mathcal{S})$ depend on the state and the joint actions of both players. The initial state distribution is denoted by $p_{0}$. At each time-step $t\leq H$, both players observe the state $s_{t}\in\mathcal{S}$ and take actions $a_{t}\in\mathcal{A}$ and $\bar{a}_{t}\in\bar{\mathcal{A}}$. The adversary is considered adaptive if its action $\bar{a}_{t}$ depends on $s_{t}$ and $a_{t}$. The protagonist receives a reward $r_{t}=R(s_{t},a_{t},\bar{a}_{t})$, while the adversary receives a negative reward $-r_{t}$.
@@ 2.1 RvS in Adversarial Environments
We extend the RvS framework by incorporating adversarial settings within the Hindsight Information Matching (HIM) paradigm [9]. In adversarial HIM, the protagonist receives a goal (denoted by $z$) and acts within an adversarial environment. The protagonist’s performance is evaluated based on how well it achieves the presented goal. Mathematically, the protagonist aims to minimize a distance metric $D(\cdot,\cdot)$ between the target goal $z$ and the information function $I(\cdot)$ applied to a trajectory $\tau$. This sub-trajectory follows a distribution $\rho$ dependent on both the protagonist’s policy $\pi_{z}=\pi(a_{t}\mid\tau_{0:t-1},s_{t},z)$ and the test-time adversarial policy $\bar{\pi}_{\textnormal{test}}$. Formally, the information distance is minimized as follows:
@@
where $\rho^{\pi_{z},\bar{\pi}_{\textnormal{test}}}$ represents the trajectory distribution obtained by executing rollouts in a Markov Game with the protagonist’s policy $\pi_{z}$ and the test-time adversarial policy $\bar{\pi}_{\textnormal{test}}$. Under HIM framework, Decision Transformer [2] uses adopt return-to-go value as the information function, while ESPER [32] employs the expected return-to-go. The well-trained RvS policy, including DT and ESPER, is derived from behavior trajectories filtered based on the condition $I(\tau)=z$:
@@
Let $\pi_{\mathcal{D}}$ and $\bar{\pi}_{\mathcal{D}}$ be the data collecting policies used to gather data for training an RvS protagonist policy $\pi$. Assume $T(s_{t+1}\mid s_{t},a_{t},{\bar{a}_{t}})=\rho^{\pi_{\mathcal{D}},\bar{\pi}_{\mathcal{D}}}(s_{t+1}\mid\tau_{0:t-1},s_{t},a_{t},{\bar{a}_{t}},I(\tau)=z)$  [^1] This condition can be readily met in environments formulated as a game with deterministic transitions, which are prevalent in many sequential games (e.g., Go and Chess [38]), as well as in frequently considered MuJoCo environments involving adversarial action noise [34, 3]. , for any goal $z$ such that $\rho^{\pi_{\mathcal{D}},\bar{\pi}_{\mathcal{D}}}(I(\tau)=z\mid s_{0})>0$, the information distance can be minimized: $\mathbb{E}_{\tau\sim\rho^{\pi_{z},\bar{\pi}_{\textnormal{test}}}}\!\left[{D(I(\tau_{t}),z)}\right]=0$ if $\bar{\pi}_{\textnormal{test}}(\bar{a}_{t}\mid\tau_{0:t-1},s_{t},a_{t})=\rho^{\pi_{\mathcal{D}},\bar{\pi}_{\mathcal{D}}}(\bar{a}_{t}\mid\tau_{0:t-1},s_{t},a_{t},I(\tau)=z)$.
@@
3 Adversarially Robust Decision Transformer
To tackle the problem of suboptimal adversarial behavior in the dataset, we propose Adversarially Robust Decision Transformer (ARDT), a worst-case-aware RvS algorithm. In ARDT, the information function is defined as the expected minimax return-to-go:
@@ 3 Adversarially Robust Decision Transformer
Intuitively, this information function represents the expected return when an adversary aims to minimize the value, while the protagonist subsequently seeks to maximize it. To train a protagonist policy as described in Eq. (3), ARDT first relabels the trajectory with the minimax returns-to-go in Eq. (4). Secondly, the protagonist policy is then trained similarly to a standard Decision Transformer (DT). This two-step overview of ARDT is provided in Figure 2, where the left block includes the minimax return estimation with expectile regression for trajectory relabeling, and the right block represents the DT training conditioned on the relabeled returns-to-go.
@@
We adopt Expectile Regression (ER) [29, 1] to approximate $\widetilde{Q}$. This choice is particularly advantageous for RL applications because it helps avoid out-of-sample estimation or the need to query actions to approximate the minimum or maximum return [16]. Specifically, the loss function for ER is a weighted mean squared error (MSE):
@@
Figure 2: Training of Adversarially Robust Decision Transformer. We adopt Expectile Regression for estimator $\widetilde{Q}$ to approximate the in-sample minimax. In the subsequent protagonist DT training, we replace the original returns-to-go with the learned values $\widetilde{Q}^{*}$ to train policy.
Algorithm 1 Adversarially Robust Decision Transformer (ARDT)
@@ 4 Experiments
In this section, we conduct experiments to examine the robustness of our algorithm, Adversarially Robust Decision Transformer (ARDT), in three settings: (i) Short-horizon sequential games, where the offline dataset has full coverage and the test-time adversary is optimal (Section 4.1), (ii) A long-horizon sequential game, Connect Four, where the offline dataset has only partial coverage and the distributional-shifted test-time adversary (Section 4.2), and (iii) The standard continuous Mujoco tasks in the adversarial setting and a population of test-time adversaries (Section 4.3). We compare our algorithm with baselines including Decision Transformer and ESPER. The implementation details are in Appendix C. Notably, since the rewards and states tokens encompass sufficient information about the adversary’s actions, all DT models are implemented not conditioned on the past adversarial tokens to reduce the computational cost.